OpenStudy (anonymous):

The equation of line AB is (y−3) = 5 (x − 4). What is the slope of a line perpendicular to line AB?

OpenStudy (anonymous):

wats sup mathmate

OpenStudy (mathmate):

First change the equation to slope-intercept form: y-3 = 5x-20 y=5x-17 The slope is 5, any line perpendicular to it will have a slope \(-\frac{1}{5}\), i.e. the negative reciprocal of this slope.

OpenStudy (anonymous):

okay i see what you did there

OpenStudy (anonymous):

hey abmon

OpenStudy (abmon98):

Hi @jamestoney Expand the brackets Y=mx+c Y-3=5x-20 Y=5x-17 m represent the gradient(slope )If Two lines with slopes m1 and m2 are perpendicular if and only if m1m2 = -1. 5*m2=-1 m2=-1/5

OpenStudy (mathmate):

If two lines are perpendicular, the product of the slopes equals -1. Or \(a_1*a_2=-1\). If you know one of the slopes, then the other slope can be found: \(a_2=-\frac{1}{a_1}\)

OpenStudy (anonymous):

ok, thanks, just needed it for future refrence

OpenStudy (mathmate):

You're welcome! :)